i need to solve this word problem with a system of 2 equations. a chemist has one solution containing 20% acid and a second solution containing 30% acid. how many liters of each solution must me combined to obtain 80 liters of a mixture that is 28% acid?
i tried the equations
L=20a+30b
80L=28(a+b)
they didnt work. please help!

Answer by checkley77(12844)

One equation will do the job.
.30x+.20(80-x)=.28*80
.30x+16-.20x=22.4
.10x=22.4=-16
.10x=6.4
x=6.4/.10
x=64 liters of 30% solution is used.
80-64=16 liters of 20% solution is used.
Proof:
.30*64+.20*16=.28*80
19.2+3.2=22.4
22.4=22.4